OpenStudy (anonymous):

Let f(x) = 2 x^2-3 x+6. f(-2 ) =

sam (.sam.):

2(-2)^2 -3(-2)+6= 20

OpenStudy (anonymous):

The average rate of change on the interval -2 <= x <= 2 is ?

OpenStudy (anonymous):

if i found f(2)=8

OpenStudy (anonymous):

and f(-2)=20

OpenStudy (anonymous):

f(x) = 2 x^2-3 x+6. f(-2) = 2* 4 - 3(-2) + 6 = 8 + 6 + 6 = 20